OpenStudy (anonymous):

((5/7) , (-2√6)/7)) Find csc t.

OpenStudy (agent0smith):

assuming those values are (cos t, sin t) \[\large \csc t =\frac{ 1 }{ \sin t } = \frac{ 1 }{ \frac{ -2 \sqrt 6 }{ 7 } } = \frac{ -7 }{ 2 \sqrt 6 } = \frac{ -7 }{ 2 \sqrt 6 }*\frac{ \sqrt 6 }{ \sqrt 6 }\]

OpenStudy (anonymous):

Thank you so much! (: It'll be \[\frac{-7\sqrt{6} }{ 12 }\] right?

OpenStudy (agent0smith):

That's right :)

OpenStudy (anonymous):

Thanks! :D

OpenStudy (agent0smith):

Welcome! :D